OpenStudy (anonymous):

A particle P is thrown vertically upwards at speed 30 ms-1 from the ground. At the same time, another particle Q, which is vertically above P, is released from rest at a height of 120 m from the ground. Determine the time taken by P to meet Q

OpenStudy (anonymous):

let p and q meet at distance x then distance covered by q=x m and by p = 120 -x m then by formula s = ut +1/2 at^2 for p s=120-x=1/2gt^2 (acceleration is gravity)-----------1 (t is time when they meet) for q s=x=30t -1/2gt^2 (u=30 and -1/2gt2 because motion opposite to gravity)----2 substituting the value of x in 2 to 1 u will get the t